OpenStudy (anonymous):

With slope -1/4 and through (8,-3)

OpenStudy (anonymous):

do what?

OpenStudy (anonymous):

y = -1/4 x + b -3 = -2 + b b = -1 equation is y = -1/4 x - 1

OpenStudy (anonymous):

they are so lazy that can't even type: find equation

OpenStudy (anonymous):

im not lazy thank you very much i just forgot to write that...

OpenStudy (anonymous):

ok